OpenStudy (anonymous):

Theo earns £20 one weekend. He gives £4 to his brother. (a) Express £4 as a fraction of £20 Give your answer in its simplest form. ..................................... (2) Theo gives £6 to his mother. (b) Express £6 as a percentage of £20 ................................. % (2) Theo spent the remaining £10 on bus fares and food. He spent £1.50 more on bus fares than on food. (c) How much did he spend on bus fares?

OpenStudy (anonymous):

is a.1/5 b.30% and c..i dunno

OpenStudy (phi):

b for bus and f for food. we can say b+f =10 £1.50 more on bus fares than on food. means that b = 1.50 more than food or b= 1.50 + f replace b with 1.5 + f in the first equation: 1.5 + f + f =10 now solve for f. can you do that ?

OpenStudy (phi):

then use b+f =10 to solve for b: b= 10-f or we could do it this way: b= 1.5+f so f= b - 1.5 replace f with b-1.5 in b+f=10 b + b - 1.50 =10 now solve for b